OpenStudy (anonymous):

I have to use sum, difference, product, or half angle formulas for this problem to find the value sin5pi/8?

OpenStudy (nayef):

I can not get your question, Can you clerify it.

zepdrix (zepdrix):

5pi/8 won't break down into addition nicely,\[\Large\rm \frac{5\pi}{8}=\frac{4\pi}{8}+\frac{\pi}{8}\]\[\Large\rm \frac{5\pi}{8}=\frac{2\pi}{8}+\frac{3\pi}{8}\]Both of those options leave us with really bad angles, 3pi/8 and the pi/8. So we'll have to try using subtraction instead.

zepdrix (zepdrix):

thinking...

zepdrix (zepdrix):

I really don't think there is a way to make that work. You instead need to use your Half-Angle Formula. Example: If trying to find \(\Large\rm \cos\left(\frac{5\pi}{8}\right)\) You would instead write it like this: \(\Large\rm \cos\left(\frac{5pi/4}{2}\right)\) And then apply your formula.

OpenStudy (anonymous):

Hmm. Question says added or substracted. So maybe just \(\huge x=\frac{5\pi}{8}\) and y=0 or viceversa

OpenStudy (anonymous):

so x+y=5pi/8 and x-y=5pi/8

OpenStudy (anonymous):

It could be something else other than addition and subtraction. I have to use sum, difference, product, or half angle formulas

OpenStudy (anonymous):

then I dont understand your question

OpenStudy (anonymous):

I'll reword it

OpenStudy (anonymous):

think of it this way you know the sin of 5pi/4 right?

OpenStudy (anonymous):

-sqrt 2/2 I'm about to get on the bus for school so I wont be able to reply quickly after this

OpenStudy (anonymous):

right 5pi/8 devided by two is 5pi/4 I hope this is enough hint

OpenStudy (anonymous):

sery, oposite :) 5pi/4 devided by 2 is 5pi/8